# Generating Chain-of-Thoughts with a Direct Pairwise-Comparison Approach to Searching for the Most Promising Intermediate Thought

### v1

@@ Abstract
To improve the ability of the large language model (LLMs) to handle complex reasoning problems, chain-of-thoughts (CoT) methods were proposed to guide LLMs to reason step-by-step, facilitating problem solving from simple to complex tasks. State-of-the-art approaches for generating such a chain involve interactive collaboration, where the learner generates candidate intermediate thoughts, evaluated by the LLM, guiding the generation of subsequent thoughts. However, a widespread yet understudied problem is that the evaluation from the LLM is typically noisy and unreliable, potentially misleading the generation process in selecting promising intermediate thoughts. In this paper, motivated by Vapnik’s principle, we propose a novel comparison-based CoT generation algorithm that directly identifies the most promising thoughts with the noisy feedback from the LLM. In each round, we randomly pair intermediate thoughts and directly prompt the LLM to select the more promising one from each pair, allowing us to identify the most promising thoughts through an iterative process. To further model the noise in the comparison, we resort to the techniques of ensemble and dueling bandits and propose two variants of the proposed algorithm. Experiments on three real-world mathematical and reasoning tasks demonstrate the effectiveness of our proposed algorithm and verify the rationale of the direct pairwise comparison.
@@ 1 Introduction
There is a fruitful line of work that considers the CoT generation problem. The pioneering work uses manual design prompts to let the LLM generate a CoT by itself (Wei et al., 2022; Wang et al., 2022). This line of research was recently extended by the score-based tree-of-thoughts (S-ToT) approaches (Yao et al., 2023; Long, 2023), where the CoT generation is framed as an interactive process with the algorithm and the LLM. These approaches generate a set of candidate intermediate thoughts each round and ask the LLM to score them and select the most promising ones. The next thoughts are then generated based on these selected ones, creating a tree-like data structure. A search algorithm, such as deep-first search, is used to identify the most promising CoT in the tree (see the illustration in Figure 2).
@@
Several preliminary approaches have been proposed to mitigate such noise in the LLM feedback, including estimating uncertainty from the semantic aspect (Kuhn et al., 2022) or ensembling multiple thoughts (Wang et al., 2022). However, getting an accurate point-wise estimate for each intermediate thought could be resource-intensive, requiring the construction of an additional model (Paul et al., 2023) or multiple queries (Wang et al., 2022): see also Figure 4 and Table 4 in our experiments. Fortunately, in the context of CoT generation, our focus is on identifying the most promising chain. Motivated by Vapnik’s principle (Vapnik, 1991), we do not need to solve a more general and difficult problem as an intermediate step, i.e. to estimate an accurate point-wise score for each intermediate thought. Meanwhile, we argue that for LLMs, comparing two thoughts simultaneously provides a more robust evaluation than assigning individual scores because it leverages more information. This argument is well established in human cognition, as seen in mathematical problems, where it is often more feasible to approximate which thought is better by comparison than by considering and evaluating them separately. We also observe this phenomenon in the experiments on the Sudoku task, as shown in Figure 1 (we leave the details to the Experiments Section), where the LLM successfully identifies the better option given two intermediate Sudoku solutions, but struggles to assign the correct value to intermediate thoughts individually. Therefore, we can use LLMs to generate a more reliable evaluation when we perform the pairwise comparison of two intermediate thoughts and directly find the most promising CoT.
@@
Based on the above insights, to alleviate the noise in the LLM feedback and to find the most promising intermediate thoughts in each round, we propose a pairwise comparison-based algorithm for CoT generation. In each round, we randomly pair all the intermediate thoughts and directly ask the LLM to compare and select the more promising one from each pair, keeping the selected one and discarding the other. Then we repeat this procedure so that we get a small set of most promising intermediate thoughts, and subsequently, we generate the next thoughts based on these selected ones. This mechanism allows us to use a direct pairwise comparison to identify the promising thoughts with a more robust evaluation. We also propose to include previous thoughts in the tree structure for comparison to mitigate the noisy nature of LLM’s feedback. Taking these two points into account, we frame the problem as an iterative process and propose a general CoT generation algorithm called comparison-based tree-of-thoughts (C-ToT). To further model the noise in the comparison, we resort to the techniques of ensemble and best-arm identification with dueling feedback (Falahatgar et al., 2017) and propose two variants of the proposed C-ToT algorithm. Through experiments on three real-world reasoning problems, we demonstrate the effectiveness of our proposed approaches and verify the rationale of the direct pairwise comparison mechanism. Our main contributions can be summarized as follows:
@@
    Motivated by Vapnik’s principle, we propose a direct pairwise-comparison approach for CoT generation that exploits noisy feedback from the LLMs.
@@ 2 Related Work
CoT Generation. Generating appropriate CoT for LLMs to enhance their inference power is a critical problem in real-world applications. Previous work has explored task-specific methods for identifying the CoT, including creating semantic graphs (Xu et al., 2021), refining the model through human-annotated CoT (Cobbe et al., 2021), or learning an additional extractor using heuristic-driven pseudo CoT (Chen et al., 2019). Different from these approaches, the LLM-based CoT generation is used directly during inference, coupling the generation process with an LLM. In these approaches, the LLM guides the CoT generation, eliminating the need for additional training.
@@
Self-Reflection. Another line of work also considered using LLMs to evaluate and to improve their own predictions. Madaan et al. (2023) and Paul et al. (2023) introduced the “self-reflection” mechanism, using the LLMs to provide feedback to their generation candidates and then fine-tuning. Paul et al. (2023) updated the model to explicitly generate intermediate thoughts while interacting with a critic model that provides automated feedback on the reasoning. Similarly, Kim et al. (2023) introduced such mechanisms in solving computer operation tasks. These methods introduced a new model to provide evaluation for the intermediate thoughts, but these critical models do not always provide perfect evaluation, so they still did not consider the noisy nature of LLM’s evaluation. Furthermore, these methods need to train an additional model, while our proposed algorithms only deal with the feedback of the LLMs during inference.
Uncertainty Quantification in LLMs. Uncertainty quantification in LLMs is a recent interest that aims to evaluate the confidence of a given answer by the LLM itself. Some work considered letting the LLM provide the confidence (Northcutt et al., 2021; Kadavath et al., 2022) by retraining the model. Another line of work considered designing entropy-based measures (Kuhn et al., 2022), or generating multiple outputs to obtain an uncertainty measure (Wang et al., 2022). Although they can be included in the CoT generation, they introduce high computation costs and latency at test time, especially if we want to get an accurate score for each intermediate thought.
Figure 2: Schematic illustration of previous CoT and S-ToT approaches with our proposed C-ToT approach for CoT generation with LLMs. Each circle box represents an intermediate thought, which is a coherent sequence of language or equations that serves as an intermediate step in problem solving. In the S-ToT method, each intermediate thought is scored by the LLM (denoted by $s$ in the figure), and the searching algorithm considers the highest-scoring ones as the most promising and then generates next intermediate thoughts based on them. In the C-ToT approach, we use direct pairwise comparison with the LLM in each round to find the most promising thoughts and then generate the next thoughts. Meanwhile, we include all previous intermediate thoughts in the comparison.
@@ 3 Our Approach
In this section, we first introduce the proposed comparison-based ToT generation algorithm, which is a general framework that generates CoT with a direct pairwise comparison method to find the most promising intermediate thought. To further alleviate the noise in LLM’s comparison feedback, we consider two different noise models and propose instantiations of our framework with theoretical analysis.
3.1 CoT Generation via Direct Pair-wise Comparison
@@
We illustrate our proposed algorithm and compare it with previous approaches in Figure 2. The CoT approaches directly ask the LLM to generate a CoT. The S-ToT approaches ask the LLM to score each intermediate thought and select the highest-scoring ones to generate the next layer. Different from these methods, we propose a direct pairwise-comparison approach to find the most promising intermediate thoughts. Note that with LLMs, due to feedback noise and input limitations, we cannot do a listwise comparison that directly asks the LLM to sort all the intermediate thoughts. Let $Z$ be the set of all candidate intermediate thoughts, and we want to select the most promising $K$ thoughts from it. The comparison iterates as follows: we randomly pair thoughts from the set and select only the winner in each pair, thereby halving the size of $Z_{i}$ to $|Z_{i}|/2$, where $Z_{i}$ denotes the set in the $i$-th iteration. After at most $K\times\log_{2}|Z|$ rounds, we can identify the $K$ most promising intermediate thoughts by such direct comparison. In practice, we can do one iteration of comparison and keep the remaining $K$ thoughts in the last few rounds. For each pair, we compare thoughts $a$ and $b$ with the LLM by asking which one is better, using different prompts with $n$ times, where $n\geq 1$. We defer the implementation details to Section 4.1.
@@
We take previous unselected intermediate thoughts into comparison to explore possibly valuable but misevaluated thoughts caused by the noise in LLM’s feedback. This is because the evaluation of intermediate thoughts may not always be accurate, and the generation of the tree structure may miss valuable intermediate thoughts in previous iterations. In the seminal research of S-ToT (Yao et al., 2023; Long, 2023; Besta et al., 2023), the thought generator uses a backtracking mechanism to revisit previous thoughts when the current ones fall below a certain threshold. While this strategy aims to rescue promising thoughts, its efficiency is questionable because it may delay the exploration of previously valuable but incorrectly scored thoughts. In addition, backtracking only occurs after a thought has fallen below a manually chosen threshold, which is difficult to know in advance for different tasks.
@@
Motivated by these shortcomings, and due to the efficiency of our direct pairwise-comparison mechanism, we maintain a repository of previous intermediate thoughts and take previous unselected thoughts into comparison at each round, rather than relying on a fixed threshold. As the example shown in Figure 2, during the pairwise comparison in the second layer, we include the intermediate thought that was not selected in the first layer in the comparison process. This ensures that the algorithm has the flexibility to backtrack to previous intermediate thoughts based on the comparison results from the LLM each round.
@@
We formulate the comparison-based ToT generation as an iterative interaction between the thought generation and the LLM. Take the C-ToT illustration in Figure 2 as an example. The algorithm starts by generating intermediate thoughts #1 to #3 based on the input. Following a direct pairwise comparison, thoughts #1 and #3 are selected, leading to the generation of new intermediate thoughts, namely #4 to #8. In the second layer, thought #2, thoughts #4 and #5 (linked to #1), along with thoughts #6 to #8 (linked to #3) are compared, subsequently resulting in the selection of thought #2 and thought #7 (linked to #3), and the generation of the next intermediate thoughts.
@@
In our approach, we keep all previous intermediate thoughts to compare in each round. This may affect the operational efficiency and exceed the storage limit. To improve the efficiency, we can introduce a counter for each intermediate thought to track its comparison frequency. If the comparison count of an intermediate thought exceeds a threshold, we can remove it from the tree. Since the comparison in each round is independent of each other, we could use parallel computing to improve the efficiency of the algorithm, or exploit more efficient machine learning techniques to schedule computational resources more adaptively and efficiently (Zhou, 2023) in the future. If the depth of the tree is $T$, the total number of comparisons required is less than $\mathcal{O}(nT\log(2mK))$.
@@
The token costs of our proposed C-ToT approach and the S-ToT approaches are task-specific and generally incomparable. The C-ToT approach could discover valuable but misevaluated previous intermediate thoughts earlier than the S-ToT method. However, it may introduce more token overhead as we compare these thoughts multiple times. Therefore, we are better suited to the problem where the initial intermediate thoughts are more uncertain. We provide the token cost analysis in Appendix C.
@@ 3.2 Instantiations and Analysis
In the proposed C-ToT framework, we use a direct comparison for each pair of thoughts. Although our C-ToT method explores two sample information compared to the S-ToT methods that use only single sample information, the comparison feedback could still be inaccurate. With different assumptions for the noise model, we can use different comparison method to select the most promising thoughts.
Sub-Gaussian Noise Assumption. Suppose the LLM returns the underlying true winner in each pair with sub-Gaussian noise. Inspired by the ensemble algorithms in CoT generation (Wang et al., 2022), we can improve the robustness of the comparison feedback by setting $n>1$ in the “Standard Mode”, so that we compare the two thoughts in each pair for $n$ times and take majority voting output.
General Noisy Comparison Assumption. We consider a more general assumption of noisy comparisons, where we only assume an unknown ranking of the $M_{t}$ thoughts at round $t$. If two thoughts $a$ and $b$ are compared, thought $a$ is chosen with some unknown probability $p(a,b)$ and $b$ is chosen with probability $p(b,a)=1-p(a,b)$, where the higher-ranked one has probability $\geq 1/2$. Repeated comparisons are independent of each other.
@@
Lemma 1 (Theorem 3 in Falahatgar et al. (2017)).
@@
Suppose that the depth of the tree is $T$ and we round Knockout each round. Then, the probability of missing the $\varepsilon$-maximum promising thoughts in the $\tau$-th layer is $1-\delta^{(T-\tau+1)}$ with $\mathcal{O}(\frac{\gamma^{2}n\sum_{i=1}^{T}|Z^{i}|}{\varepsilon^{2}}\log\frac{1}{\delta})$ comparisons.
@@
Proposition 1 is directly derived from Lemma 1 by the union bound. Proposition 1 shows that, based on the assumption of the general noisy comparison model and our proposed direct pairwise-comparison approach, the valuable intermediate thoughts will not be missed, especially for the thoughts in the shallow layer, which may be more uncertain as they appear at the beginning of the ToT generation. We leave the detailed proofs to Appendix B.

### v2

# Generating Chain-of-Thoughts with a Pairwise-Comparison Approach to Searching for the Most Promising Intermediate Thought

@@ Abstract
To improve the ability of the large language model (LLMs) to tackle complex reasoning problems, chain-of-thoughts (CoT) methods were proposed to guide LLMs to reason step-by-step, enabling problem solving from simple to complex. State-of-the-art methods for generating such a chain involve interactive collaboration, where the learner generates candidate intermediate thoughts, evaluated by the LLM, guiding the generation of subsequent thoughts. However, a widespread yet understudied problem is that the evaluation from the LLM is typically noisy and unreliable, potentially misleading the generation process in selecting promising intermediate thoughts. In this paper, motivated by Vapnik’s principle, we use pairwise-comparison evaluation instead of point-wise scoring to search for promising intermediate thoughts with the noisy feedback from the LLM. In each round, we randomly pair intermediate thoughts and directly prompt the LLM to select the more promising one from each pair, allowing us to identify the most promising thoughts through an iterative process. To further alleviate the noise in the comparison, we incorporate techniques from ensemble learning and dueling bandits, proposing two variants of the algorithm. Experiments on three real-world tasks demonstrate the effectiveness of our proposed algorithm and verify the rationale of the pairwise comparison mechanism.
Machine Learning, ICML
@@ 1 Introduction
There is a fruitful line of work that considers the CoT generation problem. The pioneering work uses manual design prompts to let the LLM generate a CoT by itself (Wei et al., 2022; Wang et al., 2022). This line of research was recently extended by the score-based tree-of-thoughts (S-ToT) approaches (Yao et al., 2023; Long, 2023), where the CoT generation is framed as an interactive process with the algorithm and the LLM. These approaches generate a set of candidate intermediate thoughts each round and ask the LLM to score them and select the most promising ones. The next thoughts are then generated based on these selected ones, creating a tree-like data structure. A search algorithm, such as deep-first search, is used to identify the most promising CoT in the tree (see the detailed illustration in Figure 2).
@@
Several preliminary approaches have been proposed to mitigate such noise in the LLM feedback, including estimating uncertainty from the semantic aspect (Kuhn et al., 2022) or ensembling multiple thoughts (Wang et al., 2022). However, getting an accurate point-wise estimate for each intermediate thought could be resource-intensive, requiring the construction of an additional model (Paul et al., 2024) or multiple queries (Wang et al., 2022): see also Figure 4 and Table 6 to 8 in our experiments. Fortunately, in the context of CoT generation, our focus is on identifying the most promising chain. Motivated by Vapnik’s principle (Vapnik, 1991), we do not need to solve a more general and difficult problem as an intermediate step, i.e. to estimate an accurate point-wise score for each intermediate thought. Instead, we can focus directly on identifying the most promising one in each round. However, it is still impractical to directly vote on all intermediate thoughts to identify the most promising one by LLMs due to the input length limit and the “lost in the middle” phenomenon (Liu et al., 2024).
We argue that for LLMs, comparing two thoughts simultaneously provides a more robust evaluation than assigning individual scores. We aim to leverage the comparison of two thoughts instead of evaluating a single thought in isolation, thereby providing a feasible alternative for identifying the most promising intermediate thought. This argument is well established in human cognition, as seen in mathematical problems, where it is often more feasible to approximate which thought is better by comparison than by considering and evaluating them separately. We also observe similar phenomena that LLMs to generate a more reliable evaluation in the experiments on the Sudoku task, as shown in Figure 1, where the LLM successfully identifies the better option given two intermediate solutions, but struggles to assign the correct value to intermediate thoughts individually.
@@
Based on the above insights, we propose a pairwise comparison-based algorithm for CoT generation to alleviate the noise in the LLM feedback and to find the most promising intermediate thoughts each round. In each round, we randomly pair all the intermediate thoughts and directly ask the LLM to compare and select the more promising one from each pair, keeping the selected one and discarding the other. Then we repeat this procedure so that we get a small set of most promising intermediate thoughts, and subsequently, we generate the next thoughts based on these selected ones. This mechanism allows us to use a direct pairwise comparison to identify the promising thoughts with a more robust evaluation. We also propose to include previous thoughts in the tree structure for comparison to mitigate the noisy nature of LLM’s feedback. Taking these two points into account, we frame the problem as an iterative process and propose a general CoT generation algorithm called comparison-based tree-of-thoughts (C-ToT). To further model the noise in the comparison, we resort to the techniques of ensemble and best-arm identification with dueling feedback (Falahatgar et al., 2017) and propose two variants of the proposed C-ToT algorithm. Through experiments on three real-world reasoning problems, we demonstrate the effectiveness of our proposed approaches and verify the rationale of the pairwise comparison mechanism. Our main contributions are:
@@
    Motivated by Vapnik’s principle, we propose a pairwise-comparison based approach for CoT generation that exploits noisy feedback from the LLMs.
@@ 2 Related Work
CoT Generation. Generating appropriate CoT for LLMs to enhance their inference power is a critical problem in real-world applications. Previous work has explored task-specific training algorithms for identifying the CoT, including creating semantic graphs (Xu et al., 2021), refining the model through human-annotated CoT (Cobbe et al., 2021), or learning an additional extractor using heuristic-driven pseudo CoT (Chen et al., 2019). Different from these approaches, the LLM-based CoT generation is used directly during inference, coupling the generation process with an LLM. In these approaches, the LLM guides the CoT generation, eliminating the need for additional training.
@@
Self-Reflection. Rather than interacting with LLMs to generate a step-by-step reasoning chain, self-reflection approaches involve LLMs directly offering an initial thought chain to the query, followed by iterative refinement of the whole chain. Madaan et al. (2023) and Paul et al. (2024) introduced the “self-reflection” mechanism, using the LLMs to provide feedback to their generation candidates and then fine-tuning. Paul et al. (2024) updated the model to explicitly generate intermediate thoughts while interacting with a critic model that provides automated feedback on the reasoning. These methods introduced new models to provide evaluation for the intermediate thoughts, but these critical models still do not always provide perfect evaluation. Furthermore, for complex problems that require sequential reasoning, such as the Game of 24, where the next thought should be generated and evaluated based on previous ones, the C-ToT generation could be more appropriate.
Uncertainty Quantification in LLMs. This is a recent interest that aims to evaluate the confidence of a given answer by the LLM itself. Some work considered letting the LLM provide the confidence (Northcutt et al., 2021; Kadavath et al., 2022) by retraining the model. Another line of work considered designing entropy-based measures (Kuhn et al., 2022), or generating multiple outputs to obtain an uncertainty measure (Wang et al., 2022). Although they can be included in the CoT generation, they introduce a high computational cost during testing, particularly when obtaining an accurate score for each intermediate thought.
Figure 2: Schematic illustration of previous CoT and S-ToT approaches with our proposed C-ToT approach for CoT generation with LLMs. Each circle box represents an intermediate thought, which is a coherent sequence of language or equations that serves as an intermediate step in problem solving. In the S-ToT method, each intermediate thought is scored by the LLM (denoted by $s$ in the figure), and the searching algorithm considers the highest-scoring ones as the most promising and then generates next intermediate thoughts based on them. In the C-ToT approach, we use pairwise comparison with the LLM in each round to find the most promising intermediate thoughts and then generate the next thoughts. Meanwhile, we include all previous intermediate thoughts in the comparison.
@@ 3 Our Approach
In this section, we first introduce the proposed comparison-based ToT generation algorithm, which is a general framework that generates CoT with a pairwise comparison mechanism to find the most promising intermediate thought. To further alleviate the noise in LLM’s comparison feedback, we propose two different instantiations of our framework with theoretical analysis.
3.1 CoT Generation via Pair-wise Comparison
@@
We illustrate our proposed algorithm and compare it with previous approaches in Figure 2. The CoT approaches directly ask the LLM to generate a CoT. The S-ToT approaches ask the LLM to score each intermediate thought and select the highest-scoring ones to generate the next layer. Different from these methods, we propose a pairwise-comparison approach to searching for the most promising intermediate thoughts. Note that with LLMs, due to feedback noise and input limitations, we cannot do a listwise voting that directly asks the LLM to sort all the intermediate thoughts. Let $Z$ be the set of all candidate intermediate thoughts, and we want to select the most promising $K$ thoughts from it. The comparison iterates as follows: we randomly pair thoughts from the set and select only the winner in each pair, thereby halving the size of $Z_{i}$ to $|Z_{i}|/2$, where $Z_{i}$ denotes the set in the $i$-th iteration. After at most $K\times\log_{2}|Z|$ rounds, we can identify the $K$ most promising intermediate thoughts by such direct comparison. In practice, we can do one iteration of comparison and keep the remaining $K$ thoughts in the last few rounds. For each pair, we compare thoughts $a$ and $b$ with the LLM by asking which one is better, using different prompts with $n$ times, where $n\geq 1$. We defer the implementation details to Section 4.1.
@@
We take previous unselected intermediate thoughts into comparison to explore possibly valuable but mis-evaluated thoughts caused by the noise in feedback. This is because the evaluation of intermediate thoughts may not always be accurate, and the generation of the tree structure may miss valuable intermediate thoughts in previous iterations. In the seminal research of S-ToT (Yao et al., 2023; Long, 2023; Besta et al., 2023), the thought generator uses a backtracking mechanism to revisit previous thoughts when the current ones fall below a certain threshold. While this strategy aims to rescue promising thoughts, its efficiency is questionable because it may delay the exploration of previously valuable but incorrectly scored thoughts. In addition, backtracking only occurs after a thought has fallen below a manually chosen threshold, which is hard to know in advance.
@@
Motivated by these shortcomings and the efficiency of our pairwise-comparison mechanism, we maintain a repository of previous intermediate thoughts. At each round, we include previously unselected thoughts in the comparisons, rather than relying on a fixed threshold. As illustrated in Figure 2, during the pairwise comparison in the second layer, we include the intermediate thought that was not selected in the first layer. This mechanism ensures that the algorithm has the flexibility to revisit previous thoughts based on the comparison results from the LLM in each round.
@@
We formulate the comparison-based ToT generation as an iterative interaction between the thought generation and the LLM. Take the C-ToT illustration in Figure 2 as an example. The algorithm starts by generating intermediate thoughts #1 to #3 based on the input. Following a pairwise comparison mechanism, thoughts #1 and #3 are selected, leading to the generation of new intermediate thoughts, namely #4 to #8. In the second layer, thought #2, thoughts #4 and #5 (linked to #1), along with thoughts #6 to #8 (linked to #3) are compared, subsequently resulting in the selection of thought #2 and thought #7 (linked to #3), and the generation of the next intermediate thoughts.
@@
In our approach, we keep all previous intermediate thoughts to compare in each round. This may affect the operational efficiency and exceed the storage limit. To improve the efficiency, we can introduce a counter for each intermediate thought to track its comparison frequency. If the comparison count of an intermediate thought exceeds a threshold, we can remove it from the tree. Since the comparison in each round is independent of each other, we could use parallel computing to improve the efficiency of the algorithm, or exploit more efficient machine learning techniques to schedule computational resources more adaptively and efficiently (Zhou, 2024) in the future. If the tree depth is $T$, the total number of comparisons required is less than the order of $\mathcal{O}(nTK\log(m))$.
@@
The token costs of our proposed C-ToT approach and the S-ToT approaches are task-specific and generally incomparable. The C-ToT approach could discover valuable but misevaluated previous intermediate thoughts earlier than the S-ToT method. However, it may introduce more token overhead as we compare these thoughts multiple times. Therefore, we are better suited to the problem where the initial intermediate thoughts are more uncertain. We provide the token cost analysis in Appendix C.2.
@@ 3.2 Instantiations and Analysis
In the proposed C-ToT framework, we use a direct comparison for each pair of thoughts. Although our C-ToT method explores two sample information compared to the S-ToT methods that use only single sample information, the comparison feedback could still be inaccurate. We offer two methods to select the winning thought in a pair.
Standard. Suppose the comparison difficulty of each pair is the same. Inspired by the ensemble algorithms in CoT generation (Wang et al., 2022), we can improve the robustness of the comparison feedback by setting $n>1$ in the “Standard Mode”, so that we compare the two thoughts in each pair for $n$ times and take majority voting output.
Dueling. We consider a more general assumption of noisy comparisons, where we only assume an unknown ranking of the $M_{t}$ thoughts at round $t$. This implies that the comparison difficulty of each pair varies, requiring a different number of comparisons for each pair. If two thoughts $a$ and $b$ are compared, thought $a$ is chosen with some unknown probability $p(a,b)$ and $b$ is chosen with $p(b,a)=1-p(a,b)$, where the higher-ranked one has probability $\geq 1/2$. Repeated comparisons are independent of each other.
@@
Lemma 1 (Theorem 3 in (Falahatgar et al., 2017)).
@@
Suppose that the depth of the tree is $T$, and thoughts in the shallower layers are more promising than those in the deeper ones. Then, the probability of missing the $\varepsilon$-maximum promising thoughts in the $\tau$-th layer is $1-\delta^{\tau}$ with at most $\mathcal{O}(\frac{\gamma^{2}\sum_{i=1}^{T}|Z^{i}|}{\varepsilon^{2}}\log\frac{1}{\delta})$ comparisons required for generating the whole tree of thoughts.
@@
Proposition 1 is directly derived from Lemma 1 by the union bound. Proposition 1 shows that, under the general assumption of noisy comparisons and utilizing our proposed pairwise-comparison approach, valuable intermediate thoughts will still not be overlooked, especially for the thoughts in the shallow layer, which may be more uncertain as they appear at the beginning of the ToT generation. We leave the detailed proofs to Appendix B.